Please.... help me with this word problem from Solving Linear Equations problem solving-word problem:
Pottery sold at the original price of 10.50 but decreased in price by $3.00 each. I f all 85 pieces sold and took in $712.50, find out how many sold at original price and how many sold at reduce price.
Thank you!

Answer by palanisamy(496)

Let x potteries be sold at the original price of 10.50 and
y potteries be sold at the reduced price of 10.50-3 = 7.5
Total number of potteries x+y = 85 ...(1)
Total amount 10.50x+7.50y = 712.50 ...(2)
(1)*7.50 => 7.50x+7.50y = 85*7.50= 637.50 ...(3)
(2)-(3)=> 3x = 75
x = 25
(1)=> 25+y = 85
y = 85-25 = 60